OpenStudy (anonymous):

Will someone PLEASE HELP ME !!!!!!! Where do the parentheses go to make the number sentence true? 8 = 24 / 4 = 2

OpenStudy (amistre64):

since 8 doesnt equal 2; then there is no set of paranthesis that will make this "true"

OpenStudy (amistre64):

8 - 24/4 = 2 regardless of paranthesis since the order of operations has us divide before we add

Directrix (directrix):

I am guesing here but given + and = on the same keyboard "key," 8 = 24 / 4 = 2 was probably meant to read: 8 = 24 / 4 + 2 If so, then --> 8 = (24 / 4) + 2. That would be 8 = 6 + 2.

OpenStudy (amistre64):

the order of operations itself satisfies the condition. 8 = 24/4 + 2 needs no paranthesis to make it a true sentence.